Let's say that we are experiencing a recession. How much will the government have to decrease taxes in order to increase GDP by $360 billion if the MPC =.5?
Answer #1

Given, MPC = 0.5

Spending multiplier can be determined using the following formula

MPS 1- MPC

MPS = 1- 0.5

MPS = 2

Now, determine the tax multiplier using the following formula

Tax Multiplier = - (MPS -1) = -(2-1) = -1

To determine increase in GDP multiply multiplier by decrease in taxes

△GDP = Multiplier * DecreaseinTax

$360billion =-1 * (-Decrease nTar)

DecreaseinTax$360billion

Thus in order to increase GDP by $ 360 billion the Government has to decrease tax by the same amount that is $ 360 billion.

Please contact if having any query thank you.